Dear all,

please find below a short summary of the activities of 25th on the mountain related to the AO system.

People working at LBTO:

G. Brusa

people working in remote:

M. Xompero (phone), A. Riccardi and R. Briguglio (VNC+Skype)


March 25th

Total # of person: 4

Summary:

Beginning of the night was spent discussing some modifications required to improve the robustness of the control GUIs of the Adaptive Secondary. AS was set around 01:40 UTC and worked without problems until about 06:00 UTC, at that time a large

current on act 500 (dsp#) sent the shell into rip state. After that event it was noticed that an area around act 356 (dsp#) had a significantly larger reading than usual in the rip state. The AS was successfully set around 06:58 UTC and kept working

for the active optics commissioning until 10:14 UTC when it was decided by the telescope operator to close the dome due to high wind (20 m/s), notice that no high wind was detected either by the local anemometer nor by the forces to be applied by

the closed loop control of the shell.

Short log:

01:40 UTC Open dome, mirror set failed, from AdOptControl, had to use Recover Failure form same GUI and was successful the second attempt. It was likely a problem of status management and not adaptive secondary hardware.

02:29 UTC wind is up to 12 m/s elevation 38 deg, it does not look like the average force on the shell has changed significantly, this appears to be in contrast with original predictions about wind effects.

04:11 UTC the images on the active optics sensor looked distorted, after re-applying the flat (using load shape) the image looks much better. It is not clear why this occurred, in principle the AS is applying a fixed shape plus an astigmatism

function of the telescope elevation.

06:14 UTC the shell was found ripped and a small set of actuators is displaying readings around 40-50, the readings appear to be off. Calling Italy for help…

06:55 UTC after some struggling trying to restart processes (VNC from Italy was not working) the shell was restarted from the mountain even with the bad area of actuators.

06:58 UTC after looking at the log it is found that actuator DSP#500 caused the rip of the shell, a max force of 0.8N was detected at ~06:00 UTC: 2010-03-26 06:00:25.517928| FUNCTEMERGENCYST > CHCURRAVERAGE-0500 = -0.8. The actuator #500 is at about 4:00 hours and ring 9. The set of actuators showing a gap larger than usual are at 7:00 hours and 12th ring.

10:10 UTC the wind is picking up, we measured 15m/s gusts outside the dining area (however at the secondary level we only measure 2m/s).

10:14 UTC the dome is closed.

10:45 UTC analysis of actuator DSP 356 shows that while at the beginning of the night in the RIP state it was measuring 32um at the end of night it measures 46um! It is not clear why this change… further investigation is required.


Ciao,

Guido


-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback